reduse 16/22 to its lowest terms.
Answer by stanbon(75887)

reduce 16/22 to its lowest terms.
----
16/22 = (2*8)/(2*11) = (2/2)(8/11) = 1*(8/11) = 8/11
==================
Cheers,
Stan H.
==========